OpenStudy (anonymous):

Given the trinomial 2x^2 + 4x + 4, predict the type of solutions. Two rational solutions One rational solution Two irrational solutions Two complex solutions

OpenStudy (anonymous):

Take the discriminate \[ b^2-4ac = (4)^2-4(2)(4) = 16-32 = -16 \]The discriminate tells us about the solutions.

OpenStudy (anonymous):

Negative discriminant tells us we have complex solutions

OpenStudy (anonymous):

it says "solutions" but only allows me to choose 1

OpenStudy (anonymous):

so that means d is eliminated?

OpenStudy (anonymous):

i meant does that mean that D is the answer